# On the Scarcity of Dense Cores ( cm) in High Latitude Planck Galactic Cold Clumps

### 26 Jan 2024 v1

@@ 4.3 Star Formation at High Galactic Latitude
Low density and high virial parameter lead to a challenge for direct gravitational collapse and then star formation of HL clouds. Observationally, it is consistent with infrared cirrus which is thought to be hostile to star formation (Low et al., 1984) and the dispersed populations of pre-main-sequence stars (see review by McGehee, 2008). Recently, a clear decreasing trend of NH (1–0) and CH (1–0) detection rates with latitude is found by Xu et al. (2021). Besides, HCN (1–0) and HCO (1–0) line survey by Braine et al. (2023) reveals that HL molecular clouds have lower dense gas fractions compared to those in the Galactic plane. Theoretically, based on Jeans mass arguments, these low-density turbulent clouds have molecular gas mass lower than the turbulent Jeans mass (cf. Table 5 in Xu et al., 2021), therefore unable to fragment into dense dust cores, or protostellar embryos, which agrees with the scarcity of dense cores observed by SCUBA-2.
@@
Previous studies have reported that the virial parameters of the PGCCs in the Taurus region (Taurus clumps hereafter) are predominantly greater than 1, with a median value of approximately 9 (Meng et al., 2013). This value is considerably lower than the median virial parameter of HLPCs, which stands at about 35. The Taurus clumps covered by the SCOPE project are designated as the Taurus SCOPE subsample. Figure,3 displays the number distribution and the DCDR for the Taurus SCOPE in orange. In the same way as above, the Mann-Whitney U test gives a p-value of , thus favoring that the DCDR of the Taurus clumps is statistically larger than that of HLPCs. Interestingly, the Taurus clumps also exhibit a significantly smaller DCDR compared to the low-latitude SCOPE clumps, as evidenced by a Mann-Whitney U test p-value of . This indicates that dense cores within the Taurus clumps are relatively rarer compared to other SCOPE clumps. Consequently, the Taurus clumps occupy an intermediate position between the HLPCs and low-latitude SCOPE clumps in terms of DCDR and virial parameter. The observed trend of decreasing DCDR with increasing virial parameter further substantiates the link between core formation efficiency and the dynamic state of the gas, as previously suggested (Eden et al., 2019).
@@
HLPCs have a distance of 200 pc which is highly consistent with the radius of the Local Bubble (LB) created by supernovae (Zucker et al., 2022). The LB is reported to expand and sweep up the ambient interstellar medium into a shell that has now fragmented and collapsed into the most prominent nearby molecular clouds. Interestingly, Zucker et al. (2022) also found that the Taurus star formation region is very likely being compressed by two super bubbles: the local super bubble and the smaller Per-Tau super bubble. If so, it’s probable that the formation of dense cores can be hindered by supernova shocks in the solar neighborhood. If the HLPCs and the Taurus clumps are on the shell, the scarcity of dense cores should favor turbulence-inhibited rather than supernova-driven star formation.
@@ 5 Conclusion
We performed a JCMT SCUBA-2 archival investigation of 70 fields towards High-Latitude Planck Galactic Cold Clumps (HLPCs) to search for dense cores. The sample benefits from being representative of the total HLPC population at low column density ( cm) and covering the densest clumps at the high column density end ( cm). Using the dust reddening in 3D map, the distances of the HLPCs are estimated to be 110–410 pc with a mean value of 200(60) pc. A total of 17 SCUBA-2 sources are identified from a mean noise rms of 15 . Only one field G6.04+36.77 (L183) contains three dense Galactic cores. The other 14 unresolved sources include 12 extragalactic objects and two Galactic young stellar objects.
Compared to the low-latitude SCOPE clumps and the Taurus clumps (at a similar distance to HLPCs), the dense core detection rate (DCDR) of HLPCs is significantly lower at the high column density end ( cm). Statistical tests verify the scarcity of dense cores in HLPCs. With synthetic observations of known dense cores, the central density of any undetected dense cores is constrained to be cm, should they exist in HLPCs. The observed scarcity of dense cores aligns with the low-density turbulent environment in HLPCs, as proposed in previous far-infrared and CO line surveys. If the HLPCs and the Taurus clumps were on the shell of the Local Bubble, the scarcity of dense cores should favor turbulence-inhibited rather than supernova-driven star formation. Furthermore, the scarcity also calls for further study on the formation mechanism of L183 dense cores.
@@ Acknowledgment
This work has been supported by the National Key R&D Program of China (No. 2022YFA1603102, 2019YFA0405100), the National Science Foundation of China (12033005, 11973013), and the China Manned Space Project (CMS-CSST-2021-A09, CMS-CSST-2021-B06), and the China-Chile Joint Research Fund (CCJRF No. 2211). CCJRF is provided by Chinese Academy of Sciences South America Center for Astronomy (CASSACA) and established by National Astronomical Observatories, Chinese Academy of Sciences (NAOC) and Chilean Astronomy Society (SOCHIAS) to support China-Chile collaborations in astronomy. T.L. acknowledges the support by the international partnership program of Chinese academy of sciences through grant No.114231KYSB20200009, and Shanghai Pujiang Program 20PJ1415500. This research was carried out in part at the Jet Propulsion Laboratory, which is operated by the California Institute of Technology under a contract with the National Aeronautics and Space Administration (80NM0018D0004). This work is sponsored (in part) by the Chinese Academy of Sciences (CAS), through a grant to the CAS South America Center for Astronomy (CASSACA) in Santiago, Chile. D.J. is supported by NRC Canada and by an NSERC Discovery Grant. G.G. acknowledges support from the ANID BASAL project FB210003. E.F. acknowledges support from the European Council, under the European Community’s Seventh framework Programme, through the Advance Grant MIST (FP7/2017-2024, No. 742719). P.S. was partially supported by a Grant-in-Aid for Scientific Research (KAKENHI Number JP22H01271 and JP23H01221) of JSPS. C.W.L. is supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2019R1A2C1010851), and by the Korea Astronomy and Space Science Institute grant funded by the Korea government (MSIT) (Project No. 2023-1-84000). G.J.W. gratefully acknowledges receipt of an Emeritus Fellowship from The Leverhulme Trust. M.J. acknowledges support from the Research Council of Finland grant 348342. The work of M.G.R. is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. N.H. acknowledge support from the National Science and Technology Council (NSTC) of Taiwan with grant NSTC 111-2112-M-001-060.
This research used the facilities of the Canadian Astronomy Data Centre operated by the National Research Council of Canada with the support of the Canadian Space Agency. The James Clerk Maxwell Telescope is operated by the East Asian Observatory on behalf of The National Astronomical Observatory of Japan; Academia Sinica Institute of Astronomy and Astrophysics; the Korea Astronomy and Space Science Institute; the National Astronomical Research Institute of Thailand; Center for Astronomical Mega-Science (as well as the National Key R&D Program of China with No. 2017YFA0402700). Additional funding support is provided by the Science and Technology Facilities Council of the United Kingdom and participating universities and organizations in the United Kingdom and Canada. Additional funds for the construction of SCUBA-2 were provided by the Canada Foundation for Innovation.
@@
Note. — The HLPC fields are listed in column (1). The extracted sources are named as SMM, as listed in column (2). The equatorial coordinates of Right Accession (RA) and Declination (DEC) in Epoch J2000 are listed in columns (3) and (4). The deconvolved standard deviation along the major and minor axis are listed in columns (5), and marked as “point” if the source is unresolved as a point source. The integrated flux and peak intensity are listed in columns (6) and (7). Identifier(s) and classifications retrieved from SIMBAD are listed in columns (8) and (9).
@@ Appendix A SCUBA-2 Observation Archive
No. | Field | RA | DEC | Project ID | Scan Pattern | Offset | rms | Distance | Altitude |
@@
Note. — The serial number and the name of HLPC are listed in columns (1) and (2). The equatorial coordinates Right Accession (RA) and Declination (DEC) of the field center in Epoch J2000 are listed in columns (3) and (4). The project ID and scan pattern of the JCMT SCUBA-2 observation are listed in columns (5) and (6). The angular offset, which is defined by the angular distance from the field center to the center of the corresponding PGCC, is listed in column (7). The rms noise of the field is listed in column (8). The distance derived from dust map is listed in column (9). The altitude from the Galactic mid-plane is listed in column (10). This table is available in its entirety in machine-readable form.

### 22 Feb 2024 v2

@@ 4.3 Star Formation at High Galactic Latitude
Low density and high virial parameter lead to a challenge for direct gravitational collapse and then star formation of HL clouds. Observationally, it is consistent with infrared cirrus which is thought to be hostile to star formation (Low et al., 1984) and the dispersed populations of pre-main-sequence stars (see review by McGehee, 2008). Recently, a clear decreasing trend of NH (1–0) and CH (1–0) detection rates with latitude is found by Xu et al. (2021). Besides, HCN (1–0) and HCO (1–0) line survey by Braine et al. (2023) reveals that HL molecular clouds have lower dense gas fractions compared to those in the Galactic plane. Theoretically, based on Jeans mass arguments, these low-density turbulent clouds have molecular gas mass lower than the turbulent Jeans mass (see Table 5 in Xu et al., 2021), therefore unable to fragment into dense dust cores, or protostellar embryos, which agrees with the scarcity of dense cores observed by SCUBA-2.
@@
Previous studies have reported that the virial parameters of the PGCCs in the Taurus region (Taurus clumps hereafter) are predominantly greater than 1, with a median value of approximately 9 (Meng et al., 2013). This value is considerably lower than the median virial parameter of HLPCs, which stands at about 35. The Taurus clumps covered by the SCOPE project are designated as the Taurus SCOPE subsample. Figure 3 displays the number distribution and the DCDR for the Taurus SCOPE in orange. In the same way as above, the Mann-Whitney U test gives a -value of , thus favoring that the DCDR of the Taurus clumps is statistically larger than that of HLPCs. Interestingly, the Taurus clumps also exhibit a significantly smaller DCDR compared to the low-latitude SCOPE clumps, as evidenced by a Mann-Whitney U test p-value of . This indicates that dense cores within the Taurus clumps are relatively rarer compared to other SCOPE clumps. Consequently, the Taurus clumps occupy an intermediate position between the HLPCs and low-latitude SCOPE clumps in terms of DCDR and virial parameter. The observed trend of decreasing DCDR with increasing virial parameter further substantiates the link between core formation efficiency and the dynamic state of the gas, as previously suggested (Eden et al., 2019).
@@
HLPCs have a distance of 200 pc which is highly consistent with the radius of the Local Bubble (LB) created by supernovae (Zucker et al., 2022). The LB is reported to expand and sweep up the ambient interstellar medium into a shell that has now fragmented and collapsed into the most prominent nearby molecular clouds. Interestingly, Zucker et al. (2022) also found that the Taurus star formation region is very likely being compressed by two super bubbles: the local super bubble and the smaller Per-Tau super bubble. If so, it is probable that the formation of dense cores can be hindered by supernova shocks in the solar neighborhood. If the HLPCs and the Taurus clumps were on the shell of LB, the scarcity of dense cores should favor turbulence-inhibited rather than supernova-driven star formation.
@@ 5 Conclusion
We performed a JCMT SCUBA-2 archival investigation of 70 fields toward HLPCs to search for dense cores. The sample benefits from being representative of the total HLPC population at low column density ( cm) and covering the densest clumps at the high column density end ( cm). Using dust reddening in 3D map, the distances of the HLPCs are estimated to be 110–410 pc with a mean value of 200(60) pc. A total of 17 SCUBA-2 sources are identified from a mean noise rms of 15 . Only one field G6.04+36.77 (L183) contains three dense Galactic cores. The other 14 unresolved sources include 12 extragalactic objects and two Galactic YSOs.
Compared to the low-latitude SCOPE clumps and the Taurus clumps (at a similar distance to HLPCs), the DCDR of HLPCs is significantly lower at the high column density end ( cm). Statistical tests verify the scarcity of dense cores in HLPCs. With synthetic observations of known dense cores, the central density of any undetected dense cores is constrained to be cm, should they exist in HLPCs. The observed scarcity of dense cores aligns with the low-density turbulent environment in HLPCs, as proposed in previous far-infrared and CO line surveys. If the HLPCs and the Taurus clumps were on the shell of the Local Bubble, the scarcity of dense cores should favor turbulence-inhibited rather than supernova-driven star formation. Furthermore, the scarcity also calls for further study on the formation mechanism of L183 dense cores.
@@ Acknowledgment
This work has been supported by the National Key R&D Program of China (No. 2022YFA1603102, 2019YFA0405100), the National Science Foundation of China (12033005, 11973013), and the China Manned Space Project (CMS-CSST-2021-A09, CMS-CSST-2021-B06), and the China-Chile Joint Research Fund (CCJRF No. 2211). CCJRF is provided by Chinese Academy of Sciences South America Center for Astronomy (CASSACA) and established by National Astronomical Observatories, Chinese Academy of Sciences (NAOC) and Chilean Astronomy Society (SOCHIAS) to support China-Chile collaborations in astronomy. T.L. acknowledges the support by the international partnership program of Chinese academy of sciences through grant No.114231KYSB20200009, and Shanghai Pujiang Program 20PJ1415500. This research was carried out in part at the Jet Propulsion Laboratory, which is operated by the California Institute of Technology under a contract with the National Aeronautics and Space Administration (80NM0018D0004). This work is sponsored (in part) by the Chinese Academy of Sciences (CAS), through a grant to the CAS South America Center for Astronomy (CASSACA) in Santiago, Chile. D.J. is supported by NRC Canada and by an NSERC Discovery grant. G.G. acknowledges support from the ANID BASAL project FB210003. E.F. acknowledges support from the European Council, under the European Community’s Seventh framework Programme, through the Advance Grant MIST (FP7/2017-2024, No. 742719). P.S. was partially supported by a Grant-in-Aid for Scientific Research (KAKENHI Number JP22H01271 and JP23H01221) of JSPS. C.W.L. is supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2019R1A2C1010851), and by the Korea Astronomy and Space Science Institute grant funded by the Korea government (MSIT) (Project No. 2023-1-84000). G.J.W. gratefully acknowledges receipt of an Emeritus Fellowship from The Leverhulme Trust. M.J. acknowledges support from the Research Council of Finland grant 348342. The work of M.G.R. is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. N.H. acknowledges support from the Nation Science and Technology Council (NSTC) of Taiwan with grant NSTC 111-2112-M-001-060.SPL acknowledges the Ministry of Science and Technology of Taiwan for grant 112-2112-M-007 -011. This research used the facilities of the Canadian Astronomy Data Centre operated by the National Research Council of Canada with the support of the Canadian Space Agency. The James Clerk Maxwell Telescope is operated by the East Asian Observatory on behalf of The National Astronomical Observatory of Japan; Academia Sinica Institute of Astronomy and Astrophysics; the Korea Astronomy and Space Science Institute; the National Astronomical Research Institute of Thailand; Center for Astronomical Mega-Science (as well as the National Key R&D Program of China with No. 2017YFA0402700). Additional funding support is provided by the Science and Technology Facilities Council of the United Kingdom and participating universities and organizations in the United Kingdom and Canada. Additional funds for the construction of SCUBA-2 were provided by the Canada Foundation for Innovation.
@@
Note. — The HLPC fields are listed in column (1). The extracted sources are named as SMM, as listed in column (2). The equatorial coordinates of R.A. and Dec. in Epoch J2000 are listed in columns (3) and (4). The deconvolved standard deviation along the major and minor axis are listed in columns (5), and marked as “point” if the source is unresolved as a point source. The integrated flux and peak intensity are listed in columns (6) and (7). Identifier(s) and classifications retrieved from SIMBAD are listed in columns (8) and (9).
@@ Appendix A SCUBA-2 Observation Archive
No. | Field | RA | Dec. | Project ID | Scan Pattern | Offset | rms | Distance | Altitude |
@@
Note. — The serial number and the name of HLPC are listed in columns (1) and (2). The equatorial coordinates Right Accession (RA) and Declination (Dec.) of the field center in Epoch J2000 are listed in columns (3) and (4). The project ID and scan pattern of the JCMT SCUBA-2 observation are listed in columns (5) and (6). The angular offset, which is defined by the angular distance from the field center to the center of the corresponding PGCC, is listed in column (7). The rms noise of the field is listed in column (8). The distance derived from dust map is listed in column (9). The altitude from the Galactic mid-plane is listed in column (10). This table is available in its entirety in machine-readable form.